Find the area of the shaded region. The graph depicts the standard normal distribution of bone density scores with mean 0 and standard deviation 1.

z= -0.84

A symmetric bell-shaped curve is plotted over a horizontal scale. A vertical line runs from the scale to the curve at labeled coordinate “z equals negative 0.84,” which is to the left of the curve’s center and peak. The area under the curve to the right of the vertical line is shaded.

The area of the shaded region is ​(Round to four decimal places as​ needed.)
Answer #1

By considering the given information,

Area of shaded region is

P ( Z>=-0.84) = 1 - P ( Z <= -0.84)

= 1 - P ( Z >= 0.84) ..... Due to symmetry of normal distribution

= 1 - 0.20045

= 0.7996

Area of shaded region is 0.7996